# EVTAR: End-to-End Try on with Additional Unpaired Visual Reference

### v1

@@ 3.3 Muti-input Adaptation of EVTAR
After obtaining the unpaired person images, we train our end-to-end model by replacing the agnostic images $\bm{a}_{i}$ with the corresponding person images $\bm{p}_{i}$, which are generated by the first stage model. To enable the model to handle both masked and unmasked inputs flexibly, we randomly feed agnostic and person images with equal probability (50%) during training. Instead of training a try-on model from scratch, we adopt Flux-Kontext [46] as the foundation model, freeze the parameters of its encoder and decoder, and apply Low-Rank Adaptation (LoRA) [48] to the MM-DiT blocks for parameter-efficient fine-tuning. The training objective follows the same flow-matching loss as defined in Equation 2, where $\bm{a}_{i}$ and $\bm{p}_{i}$ are alternately used to ensure compatibility with both input types. To enable the model to extract visual information from the target garment worn by another person, an additional reference image $\bm{r}_{i}$ is provided as input with a certain probability ($25\%$). Its latent embeddings are appended to the latent feature sequence $\bm{\epsilon}_{i},\bm{a}_{i},\bm{c}_{i}]$ after the image encoder, as illustrated in Fig. 3(b).
Figure 4: Generating the tryon result with additional visual reference. We train a tryon model conditioned on cloth $\bm{c}_{i}$, reference image $\bm{r}_{i}$, and takes agnostic $\bm{a}_{i}$ or reference image $\bm{r}_{i}$ as input to fit the cloth image onto the target person.
The Flux-Kontext architecture provides a flexible framework for image editing. The input image is first converted into latent embeddings by the image encoder and then flattened into a sequence of shape $[L,D]$, where $L$ denotes the sequence length per image and $D$ is the embedding dimension. This sequence is concatenated with the Gaussian noise $\bm{\epsilon}$ along the sequence dimension. Similar to a Vision Transformer that employs positional embeddings to encode spatial information, Flux-Kontext uses an index generator to produce a positional index of shape $[L,3]$ for the image embedding sequence. These indices are subsequently transformed into Rotary Positional Embeddings (RoPE), providing additional spatial cues to the model. The first channel of the positional index indicates whether the embedding corresponds to Gaussian noise $\bm{\epsilon}_{i}$ or conditional inputs, while the second and third channels represent the horizontal and vertical positional indices, respectively, as illustrated in Fig. 4.
Based on this design, we first convert the conditional inputs into latent embeddings with a shape of $[L,D]$, denoted as $\bm{a}_{i},\bm{c}_{i},\bm{r}_{i}]$, and concatenate them with the Gaussian noise embeddings $\bm{\epsilon}_{i}$ along the sequence dimension. However, the vanilla Flux-Kontext [46] supports only a single conditional image input, which is insufficient for virtual try-on tasks that require at least the target person $\bm{p}_{i}$ and the garment $\bm{c}_{i}$. To overcome this limitation, we extend the positional index for RoPE used by the DiT blocks. Each embedding is assigned a three-channel index: the first channel, originally binary (0 for noise and 1 for condition), is expanded to values from 0 to 3 to distinguish multiple conditional inputs; the second and third channels encode horizontal and vertical positions as integers. To support multi-resolution conditions, we further scale the positional indices in the second and third channels according to the resolution ratio between the target and conditional images, ensuring that all indices remain within a consistent numerical range. The modified positional index design of the DiT blocks and the overall architecture of our EVTAR are illustrated in Fig. 4.
@@ 3.4 Virtual-Tryon Generation with Extra Visual Reference
Figure 5: The overall pipeline of generating the reference images. We first generate the appearance descriptions using Qwen2.5-VL [49], and then concatenate the appearance with the corresponding actions and outfits to construct the positive and negative prompts, as shown in (a). Subsequently, the images and the textual prompts are fed into the Editing Model, which generates photos of individuals wearing the same clothes. These results are reference images for each image–cloth pair, as shown in (b).
For humans, particularly in e-commerce shopping, we believe that relying solely on garment images and conditions extracted by external models, as done in previous methods, is insufficient to imagine the realistic visual effects of fitting a target garment onto one’s body. A garment’s style, texture, and fine design elements are more faithfully represented when worn by another person, whereas viewing the garment in isolation fails to convey these nuances. The same applies to generative virtual try-on models, as illustrated in Fig. 2. Therefore, we argue that relying on cloth images and conditions extracted by external models, as in previous methods, is insufficient to capture the realistic visual effect of fitting a target garment onto a person. By incorporating reference person images, that the target clothes worn by another person $\bm{r}_{i}$, into the virtual try-on generation process, the model can similarly get more intuitive visual guidance.
To provide the model with reference person images $\bm{r}_{i}$, we construct pairs in the form of “different persons wearing the target garment,” denoted as $\bm{c}_{i},\bm{r}_{i}]$. Such reference images $\bm{r}_{i}$ are not available in existing open-source virtual try-on datasets, similar to the person images $\bm{p}_{i}$. In this subsection, we introduce our data engine, which synthesizes the required reference images, enriching current virtual try-on datasets with supplementary references and enabling our model to be trained with additional visual guidance.
Existing open-source datasets, including VITON-HD [4], DressCode [26], and IGPairs [29], lack unpaired reference images $\bm{r}_{i}$ showing the different persons wearing the target garment, thereby constraining the model’s training. We employ generative models to synthesize such reference images to overcome this limitation. For the generated reference data $\bm{r}_{i}$ to provide accurate and intuitive visual guidance, we argue that they should meet the following requirements:
  1. 1.
@@
  2. 2.
    Introduce diversity in person identity. The person wearing the target garment in the reference image should not be identical to the target person wearing the same garment. Otherwise, the model may overfit to the target image. This diversity can be achieved by altering hairstyle, hair/skin color, body pose, or facial expressions.
  3. 3.
    Vary the non-target garments. While the target garment remains unchanged, other garments should be modified. For example, if the target garment is an upper-body item, the reference image should retain the same upper-body garment but alter the lower-body clothing, shoes, or accessories.
@@
As shown in the pipeline, we employ Flux-Kontext [46] to generate reference images from the target descriptions, leveraging its strong capability to maintain consistency with the input image (1). To address requirement (2), we describe the target image and then intentionally provide an alternative description that ensures the reference image differs from the target input. For this purpose, we use Qwen2.5-VL [49] to generate detailed descriptions of the model’s appearance, and instruct it to produce variants that do not resemble the original model. Finally, to fulfill requirement (3), we curate a list of garment descriptions representing diverse non-target clothing items. The outlook, action, and outfit descriptions are concatenated as the positive prompt. In contrast, the original image description is used as the negative prompt, and both are fed into the generative model to synthesize reference images.
We supplement existing virtual try-on benchmarks including VITON-HD [5], DressCode [26], ViViD [27], FashionTryOn [28] and IGPairs [29] by generating corresponding reference pairs for each target garment image $\bm{c}_{i}$, forming data pairs $\bm{c}_{i},\bm{r}_{i}]$ that are used for both model training and evaluation. Some open-source datasets, such as FashionTryOn [28] and IGPairs [29], contain numerous duplicated or low-quality samples. We filter out redundant or poor-quality images to ensure overall data quality before generating the final reference data using our data engine. Finally, we enrich open-source virtual try-on datasets with additional visual references $\bm{r}_{i}$ and person images $\bm{p}_{i}$, and combine them to form our own dataset, named Virtual Fitting with Reference (VFR), for training our EVTAR model.
@@ 4 Experiments
We evaluate the effectiveness of our proposed method on two public benchmarks, DressCode [26] and VITON-HD [4], both containing images with a resolution of $1024\times 768$. The VITON-HD dataset comprises 13,670 upper-body image pairs of women, split into 11,647 pairs for training and 2,032 pairs for testing. The DressCode dataset includes three subsets—upper body, lower body, and dresses—with 48,392 training and 5,400 testing pairs. Since DressCode does not provide wrapped cloth masks or agnostic images $\bm{a}_{i}$, we generate them using the mask generation tool from CatVTON [14]. During training, we fine-tune our model on the flux-kontext backbone using the Low-Rank Adaptation (LoRA) technique, with a rank of $64$ and $\alpha=128$, optimized by AdamW. For quantitative evaluation, all generated images are resized to $512\times 384$ to ensure a fair comparison with previous methods. In a single-dataset experiment, the model is trained independently for 20,000 steps on VITON-HD and 48,000 on DressCode with a batch size of 128, using 8 NVIDIA H100 GPUs. To further enhance generalization and robustness to in-the-wild inputs, we train an additional model on the mixed VFR dataset introduced in Sec. 3.4. We report cross-dataset evaluation results of our EVTAR model at a resolution of $1024\times 768$ on both VITON-HD and DressCode, and demonstrate its in-the-wild performance using images captured by ourselves and collected from the internet.
@@ 4.1 Quantative result
We evaluate the numerical results of our virtual try-on model on the VITON and DressCode datasets, distinguishing between paired and unpaired try-on settings. For paired try-on settings with ground truth in test datasets, we utilize four widely used metrics to assess the similarity between the synthesized images and their corresponding authentic images: the Structural Similarity Index (SSIM) [50], Learned Perceptual Image Patch Similarity (LPIPS) [51], Fréchet Inception Distance (FID)[52], and Kernel Inception Distance (KID)[53]. For unpaired settings, where we measure the distributional similarity between the synthesized and real samples, we specifically rely on the Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). As shown in Table1.
Table 1: Quantitative results on the VITON-HD dataset [4]. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous work. The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. In the Input columns, a “✓” symbol indicates that the model requires the corresponding input, while “–” means it does not.
@@
Mask Pose Prompt
CAT-DM [54] 0.080 0.877 5.60 0.83 8.93 1.37
IDM-VTON [12] 0.102 0.870 6.29
OOTDiffusion[8] 0.071 0.878 8.81 0.82
CatVTON [14] 0.057 0.870 5.43 0.41 9.02 1.09
CatVT2ON [15] 0.057 0.890 8.10 2.25 11.22 2.99
OmniVTON [16] 0.145 0.832 7.76 9.62
PromptDresser pose [13] 0.097 0.878 9.07 1.16
PromptDresser [13] 0.112 0.869 8.54 0.67
EVTAR(Ours) 0.057 0.873 5.45 0.82 8.58 1.06
EVTAR+R(Ours) 0.049 0.879 4.69 0.68 8.43 0.91
EVTAR/MF (Ours) 0.061 0.866 5.98 1.04 8.40 0.81
EVTAR+R/MF(Ours) 0.053 0.872 5.11 0.82 8.32 0.78
Our method (EVTAR) consistently performs better than prior baselines, demonstrating higher try-on fidelity and strong alignment with the target person’s pose. With the addition of reference images (“+R”), the quality and detail consistency of the try-on results is further improved compared with the results without reference images, establishing new state-of-the-art results across multiple metrics. Notably, even in the mask-free setting—without agnostic masks or auxiliary inputs—our method maintains garment style correctness and pose consistency, while reaching accuracy on par with or superior to baseline methods, highlighting its robustness and practicality.
Table 2: Quantitative results on the DressCode dataset [26]. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous work. The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively.
Method All
@@
IDM-VTON  [4] 0.062 0.920 8.64 0.90
OOTDiffusion  [8] 0.045 0.927 4.20 0.37
CatVTON  [14] 0.046 0.892 3.99 0.82 6.14 1.40
CatVT2ON  [15] 0.037 0.922 5.72 2.34 8.63 3.84
OmniVTON  [16] 0.119 0.865 5.34 6.45
EVTAR(Ours) 0.037 0.912 3.48 1.20 5.31 1.36
EVTAR+R(Ours) 0.031 0.918 2.94 0.95 5.07 1.15
EVTAR/MF (Ours) 0.041 0.901 3.84 1.33 5.00 1.17
EVTAR+R/MF(Ours) 0.035 0.906 3.34 1.15 5.02 1.28
Table 3: Quantitative results on different subsets of the Dress Code dataset [26]. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous work. The symbol “+R” denotes results with reference image input, and “MF” indicates mask-free input images. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively.
@@
CAT-DM [54] 9.85 2.38 12.62 1.89 10.25 1.81 14.83 2.82 10.71 2.02 14.30 3.36
OOTDiffusion  [8] 11.03 0.29 9.72 0.64 10.65 0.54
PromptDresser [13] 11.00 0.74 12.55 1.46 11.09 1.10
EVTAR(Ours) 7.62 1.10 11.13 0.98 7.60 1.38 13.07 2.11 7.32 1.30 11.56 1.98
EVTAR+R(Ours) 6.39 0.85 11.08 0.87 6.61 1.05 12.56 1.67 6.09 1.16 11.16 1.72
EVTAR/MF(Ours) 8.37 1.43 11.20 1.11 8.79 1.51 12.50 1.83 7.36 1.33 10.73 1.41
EVTAR+R/MF(Ours) 7.20 1.18 11.53 1.12 7.85 1.21 12.74 2.05 6.24 1.20 10.05 1.30
Table 2 and Table 3 summarizes the quantitative evaluation on the DressCode dataset. Our method (EVTAR) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality.

### v2

@@ 3.3 Multi-input Adaptation of RefTON
After obtaining the unpaired person images, we train our person-to-person model by replacing the agnostic images $\bm{a}_{i}$ with the corresponding person images $\bm{p}_{i}$, which the first stage model generates. To enable the model to handle both masked and unmasked inputs flexibly, we randomly feed agnostic and person images with equal probability (50%) during training. Instead of training a try-on model from scratch, we adopt Flux-Kontext [labs2025flux1kontextflowmatching] as the foundation model, freeze the parameters of its encoder and decoder, and apply Low-Rank Adaptation (LoRA) [hu2022lora] to the MM-DiT blocks for parameter-efficient fine-tuning. The training objective follows the same flow-matching loss as defined in Equation 2, where $\bm{a}_{i}$ and $\bm{p}_{i}$ are alternately used to ensure compatibility with both input types. To enable the model to extract visual information from the target garment worn by another person, an additional reference image $\bm{r}_{i}$ is provided as input with a certain probability ($25\%$). Its latent embeddings are appended to the latent feature sequence $\bm{\epsilon}_{i},\bm{a}_{i},\bm{c}_{i}]$ after the image encoder, as illustrated in Fig. 4.
Some prior work [guo2025any2anytryon] leverages adaptive position embeddings for virtual try-on. Similarly, Flux-Kontext employs an index generator to produce a position index of shape $[L,3]$ for the image embedding sequence, which is then transformed into Rotary position Embeddings (RoPE) to provide spatial cues. The first channel distinguishes Gaussian noise $\bm{\epsilon}_{i}$ from conditional embeddings, while the second and third encode horizontal and vertical coordinates (Fig. 4).
Building on this design, our model converts conditional inputs into latent feature after Auto-Encoder and flatten into $L$ $D$-dimensional latent embeddings, denoted as $\bm{a}_{i},\bm{c}_{i},\bm{r}_{i}]$, and concatenates them with Gaussian noise embeddings $\bm{\epsilon}_{i}$ along the sequence dimension. This formulation allows conditional inputs of multi-resolutions to be concatenated—unlike [guo2025any2anytryon], which concatenates noisy and conditional inputs directly in pixel space, requiring them to share the same image size. Each embedding is assigned a three-channel position index: the first channel is extended from a binary flag to four discrete values to distinguish multiple conditions, and the second and third channels store integer spatial coordinates. To handle multi-resolution conditions, we further rescale these position indices by the resolution ratio between the target and conditional images, ensuring numerical consistency across resolutions. The modified DiT position index design and the overall RefTON architecture are illustrated in Fig. 4.
Figure 5: The overall pipeline of generating the reference images. We first generate the appearance descriptions using Qwen2.5-VL [bai2025qwen2], and then concatenate the appearance with the corresponding actions and outfits to construct the positive and negative prompts, as shown in (a). Subsequently, the images and the textual prompts are fed into the Editing Model, which generates photos of individuals wearing the same clothes. These results are reference images for each image–cloth pair, as shown in (b).
@@ 3.4 Virtual-Tryon Generation with Extra Visual Reference
For humans, particularly in e-commerce shopping, we believe that relying solely on garment images and conditions extracted by external models, as done in previous methods, is insufficient to imagine the realistic visual effects of fitting a target garment onto one’s body. A garment’s style, texture, and fine design elements are more faithfully represented when worn by another person, whereas viewing the garment in isolation fails to convey these nuances. The same applies to generative virtual try-on models, as illustrated in Fig. 2. Therefore, we argue that relying on cloth images and conditions extracted by external models, as in previous methods, is insufficient to capture the realistic visual effect of fitting a target garment onto a person. By incorporating reference person images, that the target clothes worn by a different person $\bm{r}_{i}$, into the virtual try-on generation process, the model can similarly get more direct and intuitive visual guidance.
To provide the model with reference person images $\bm{r}_{i}$, we construct pairs in the form of “different persons wearing the target garment,” denoted as $\bm{c}_{i},\bm{r}_{i}]$. Such reference images $\bm{r}_{i}$ are not available in existing open-source virtual try-on datasets, similar to the person images $\bm{p}_{i}$. In this subsection, we introduce our reference data generation pipeline, which synthesizes the required reference images, enriching current virtual try-on datasets with supplementary references and enabling our model to be trained with additional visual guidance.
Existing open-source datasets, including VITON-HD [choi2021viton], DressCode [morelli2022dress], and IGPairs [shen2025imagdressing], lack unpaired reference images $\bm{r}_{i}$ showing the different persons wearing the target garment, thereby constraining the model’s training. We employ generative models to synthesize such reference images to overcome this limitation. For the generated reference data $\bm{r}_{i}$ to provide accurate and intuitive visual guidance, we argue that they should meet the following requirements:
@@
Introduce diversity in person’s appearance. The person wearing the target cloth in the reference image should unlike the target person wearing the same garment. Otherwise, the model may learn shortcut and overfit to the target image. This diversity can be achieved by altering hairstyle, hair/skin color, body pose, or facial expressions.
Vary the non-target garments to provide outfit diversity. While the target garment remains unchanged, other garments should be modified. For example, if the target garment is an upper-body item, the reference image should retain the same upper-body garment but alter the lower-body clothing, shoes, or accessories.
@@
As shown in the pipeline, we employ Flux-Kontext [labs2025flux1kontextflowmatching] to generate reference images from the target descriptions, leveraging its strong capability to maintain consistency with the input image (1). To address requirement (2), we describe the target image and then intentionally provide an alternative description that ensures the reference image differs from the target input. For this purpose, we use Qwen2.5-VL [bai2025qwen2] to generate detailed descriptions of the model’s appearance, and instruct it to produce variants that do not resemble the original model. Finally, to fulfill requirement (3), we curate a list of garment descriptions representing diverse non-target clothing items. The outlook, action, and outfit descriptions are concatenated as the positive prompt. In contrast, the original image description is used as the negative prompt, and both are fed into the generative model to synthesize reference images.
We supplement existing virtual try-on benchmarks including VITON-HD [han2018viton], DressCode [morelli2022dress], ViViD [fang2024vivid], FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing] by generating corresponding reference pairs for each target garment image $\bm{c}_{i}$, forming data pairs $\bm{c}_{i},\bm{r}_{i}]$ that are used for both model training and evaluation.Some open-source datasets, such as FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing], contain numerous duplicated or low-quality samples. We compare the CLIP features of images to filter out redundant samples and employ Qwen2.5-VL to identify distorted or unclear images, as well as images where the person faces away from the camera, ensuring overall data quality before generating the final reference set via our data generation pipeline. Finally, we enrich open-source virtual try-on datasets with additional visual references $\bm{r}_{i}$ and person images $\bm{p}_{i}$, and combine them to form our own dataset, named Virtual Fitting with Reference (VFR), for training our RefTON model. The detailed data collection, filtering procedures, and sample’s visualization are provided in the Appendix.
@@ 4 Experiments
We evaluate the effectiveness of our proposed method on two public benchmarks, DressCode [morelli2022dress] and VITON-HD [choi2021viton], both containing images with a resolution of $1024\times 768$. The VITON-HD dataset comprises 13,670 upper-body image pairs of women, split into 11,647 pairs for training and 2,032 pairs for testing. The DressCode dataset includes three subsets—upper body, lower body, and dresses—with 48,392 training and 5,400 testing pairs. Since DressCode does not provide wrapped cloth masks or agnostic images $\bm{a}_{i}$, we generate them using the mask generation tool from CatVTON [chong2024catvton]. During training, we fine-tune our model on the flux-kontext backbone using the Low-Rank Adaptation (LoRA) technique, with a rank of $64$ and $\alpha=128$, optimized by AdamW. For quantitative evaluation, all generated images are resized to $512\times 384$ to ensure a fair comparison with previous methods. In a single-dataset experiment, the model is trained independently for 20,000 steps on VITON-HD and 48,000 on DressCode with a batch size of 128, using 8 NVIDIA H100 GPUs. To further enhance generalization and robustness to in-the-wild inputs, we train an additional model on the mixed VFR dataset introduced in Sec. 3.4. We report cross-dataset evaluation results of our RefVTON model at a resolution of $1024\times 768$ on both VITON-HD and DressCode, and demonstrate its in-the-wild performance using images captured by ourselves and collected from the internet.
@@ 4.1 Quantative result
We evaluate the numerical results of our virtual try-on model on the VITON and DressCode datasets, distinguishing between paired and unpaired try-on settings. For paired try-on settings with ground truth in test datasets, we utilize four widely used metrics to assess the similarity between the synthesized images and their corresponding authentic images: the Structural Similarity Index (SSIM) [nilsson2020understanding], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Fréchet Inception Distance (FID)[Seitzer2020FID], and Kernel Inception Distance (KID)[binkowski2018demystifying]. For unpaired settings, where we measure the distributional similarity between the synthesized and real samples, we specifically rely on the Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). As shown in Table1.
Our method (RefVTON) consistently performs better than prior baselines, demonstrating higher try-on fidelity and strong alignment with the target person’s pose. With the addition of reference images (“+R”), the quality and detail consistency of the try-on results is further improved compared with the results without reference images, establishing new state-of-the-art results across multiple metrics. Notably, even in the mask-free setting—without agnostic masks or auxiliary inputs—our method maintains garment style correctness and pose consistency, while reaching accuracy on par with or superior to baseline methods, highlighting its robustness and practicality.
Table 1: Quantitative comparison across VITON-HD [choi2021viton] and DressCode [morelli2022dress]. The best and second best results are shown in bold and underline. “+R” denotes the use of reference images, and “MF” indicates mask-free inputs. Missing values are shown as “–”. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. Unless otherwise specified, the same notations carry the same meanings throughout the figures and tables in this paper.
Method Input VITON-HD DressCode
@@
CAT-DM [zeng2024cat] 0.080 0.877 5.60 0.83 8.93 1.37
IDM-VTON [choi2024improving] 0.102 0.870 6.29 0.062 0.920 8.64 0.90
OOTDiffusion [xu2025ootdiffusion] 0.071 0.878 8.81 0.82 0.045 0.927 4.20 0.37
CatVTON [chong2024catvton] 0.057 0.870 5.43 0.41 9.02 1.09 0.046 0.892 3.99 0.82 6.14 1.40
CatVT2ON [chong2025catv2ton] 0.057 0.890 8.10 2.25 11.22 2.99 0.037 0.922 5.72 2.34 8.63 3.84
OmniVTON [yang2025omnivton] 0.145 0.832 7.76 9.62 0.119 0.865 5.34 6.45
@@
PromptDresser [kim2024promptdresser] 0.112 0.869 8.54 0.67
RefTON (Ours) 0.057 0.873 5.45 0.82 8.58 1.06 0.037 0.912 3.48 1.20 5.31 1.36
RefTON+R (Ours) 0.049 0.879 4.69 0.68 8.43 0.91 0.031 0.918 2.94 0.95 5.07 1.15
Mask-Free setting
CatVTON(MF) [chong2024catvton] 0.061 0.870 5.89 0.51 9.29 1.17 0.045 0.902 4.78 1.30 7.40 2.62
Any2AnyTryon [guo2025any2anytryon] 0.088 0.839 6.93 0.74 8.97 0.98
TryOffDiff [velioglu2025enhancing] 11.9 2.60 7.90 2.70
RefTON/MF (Ours) 0.061 0.866 5.98 1.04 8.40 0.81 0.041 0.901 3.84 1.33 5.00 1.17
RefTON+R/MF(Ours) 0.053 0.872 5.11 0.82 8.32 0.78 0.035 0.906 3.34 1.15 5.02 1.28
Table 2: Quantitative results on three subsets of the DressCode dataset [morelli2022dress]: upper body, lower body, and dresses. The best and second-best results are highlighted in bold and underline, respectively. The symbol “*” denotes results reported in prior work, while “+R” indicates results with reference image input, and “MF” refers to mask-free input images. The subscripts $p$ and $s$ denote specific evaluation metrics for precision and recall, respectively.
@@
CAT-DM [zeng2024cat] 9.85 2.38 12.62 1.89 10.25 1.81 14.83 2.82 10.71 2.02 14.30 3.36
OOTDiffusion  [xu2025ootdiffusion] 11.03 0.29 9.72 0.64 10.65 0.54
PromptDresser [kim2024promptdresser] 11.00 0.74 12.55 1.46 11.09 1.10
RefTON(Ours) 7.62 1.10 11.13 0.98 7.60 1.38 13.07 2.11 7.32 1.30 11.56 1.98
RefTON+R(Ours) 6.39 0.85 11.08 0.87 6.61 1.05 12.56 1.67 6.09 1.16 11.16 1.72
RefTON/MF(Ours) 8.37 1.43 11.20 1.11 8.79 1.51 12.50 1.83 7.36 1.33 10.73 1.41
RefTON+R/MF(Ours) 7.20 1.18 11.53 1.12 7.85 1.21 12.74 2.05 6.24 1.20 10.05 1.30
Table 1 and Table 2 summarizes the quantitative evaluation on the DressCode dataset. Our method (RefVTON) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality.